# Identification of Semiparametric Panel Multinomial

### v2

@@ Abstract
This paper proposes a robust method for semiparametric identification and estimation in panel multinomial choice models, where we allow for infinite-dimensional fixed effects that enter into consumer utilities in an additively nonseparable way, thus incorporating rich forms of unobserved heterogeneity. Our identification strategy exploits multivariate monotonicity in parametric indexes, and uses the logical contraposition of an intertemporal inequality on choice probabilities to obtain identifying restrictions. We provide a consistent estimation procedure, and demonstrate the practical advantages of our method with Monte Carlo simulations and an empirical illustration on popcorn sales with the Nielsen data.
@@ 1 Introduction
This paper proposes a method for semiparametric identification and estimation in panel multinomial choice models, where we allow for infinite-dimensional fixed effects that enter into consumer utilities in an additively nonseparable manner. The proposed method also applies much more widely beyond panel multinomial choice models, and can be easily adapted to a wide range of models characterized by multi-index single-crossing conditions, which we introduce later in this paper.
@@
$y_{ijt}=\mathbf{\mathbbm{1}}\left\{u\left(X_{ijt}^{{}^{\prime}}\beta_{0},\,A_{ij},\,\epsilon_{ijt}\right)\geq\max_{k\in\left\{1,...,J\right\}}u\left(X_{ikt}^{{}^{\prime}}\beta_{0},\,A_{ik},\,\epsilon_{ikt}\right)\right\}$
@@
where agent $\mathit{i}$’s utility from a candidate product $j$ at time $t$, represented by $u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)$, is taken to be a function of three components. The first is a linear index $X_{ijt}^{{}^{\prime}}\beta_{0}$ of observable characteristics $X_{ijt}$, which contains a finite-dimensional parameter of interest $\beta_{0}$ we will identify and estimate. The second term $A_{ij}$ is an infinite-dimensional fixed effect that can be heterogeneous across each agent-product combination. We emphasize that $X_{ijt}$ and $A_{ij}$ can be arbitrarily dependent. The last term $\epsilon_{ijt}$ is an idiosyncratic time-varying error term of arbitrary dimensions. The three components are then aggregated by an unknown utility function $u$ in an additively nonseparable way, with the only restriction being that each agent’s utility $u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)$ is increasing in its first argument. Each agent then chooses a certain product in a given time period, represented by $y_{ijt}=1$, if and only if this product gives him the highest utility among all available products.
The infinite-dimensionality of the terms $u$, $A_{ij}$ and $\epsilon_{ij}$ and the additive nonseparability in their interactions jointly produce rich forms of unobserved heterogeneity. Across each agent-product combination $ij$, we are effectively allowing for nonparametric variations in agent utilities as functions of the index $X_{ijt}^{{}^{\prime}}\beta_{0}$, which proxies for the effects of complicated unobserved factors that influence choice behavior, such as brand loyalty, subtle flavors, and unique styles of products. Moreover, unrestricted heterogeneity in the distribution of the error term $\epsilon_{ijt}$ is accommodated.
@@
The generality of our setup encompasses many semiparametric (or parametric) panel multinomial choice models with scalar-valued fixed effects, scalar-valued error terms and various degrees of additive separability in the related literature, including the following standard formulation:
@@
$y_{ijt}=\mathbf{\mathbbm{1}}\left\{X_{ijt}^{{}^{\prime}}\beta_{0}+A_{ij}+\epsilon_{ijt}\geq\max_{k\in\left\{1,...,J\right\}}\left(X_{ikt}^{{}^{\prime}}\beta_{0}+A_{ik}+\epsilon_{ikt}\right)\right\}.$
@@
Relatively speaking, in this paper we are able to accommodate the infinite dimensionality of unobserved heterogeneity and the lack of additive separability in agent utility functions, under a standard time homogeneity assumption on the idiosyncratic error term that is widely adopted in the related literature.
@@
Our key identification strategy exploits the standard notion of multivariate monotonicity in its contrapositive form. The idea is very simple and intuitive: whenever we observe a strict increase in the choice probabilities of a specific product from one period to another, by logical contraposition it cannot be possible that this product becomes worse while all other products become better over the two periods. Hence, by taking the logical contraposition of a carefully constructed inequality on conditional choice probabilities, we can obtain an identifying restriction on the index values free of all infinite-dimensional nuisance parameters, with which we construct a population criterion function.
@@
Based on our identification result, we provide consistent two-step set (or point) estimators, together with a computational algorithm adapted to the technical challenges of our framework. The first stage takes the form of a standard nonparametric regression, where we estimate a collection of intertemporal differences in conditional choice probabilities. In the second stage, we numerically minimize our sample criterion function with the first-stage estimates plugged in. A highlight of our computation procedure is the adoption of a spherical-coordinate reparameterization of our criterion functions in terms of angles, which enables us to exploit a combination of topological, geometric and computational advantages. A simulation study is conducted to analyze the finite-sample performance of our method and the adequacy of our computational procedure for practical implementation.
@@
We also provide an empirical illustration of our procedure, where we use the Nielsen data on popcorn sales in the United States to explore the effects of marketing promotion effects. The results show that our procedure produces estimates that conform well with economic intuition. For example, we find that special in-store displays boost sales not only through a direct promotion effect but also through the attenuation of consumer price sensitivity, a result that cannot be produced by other methods based on additive separability. Intuitively, marketing managers are more likely to promote products that they know consumers are more price and promotion sensitive to. Hence, the average effective price sensitivity of promoted products tend to be larger than those not promoted due to the selection effect. Given the nonadditive nature of such selection effects, estimators based on additive separability will be biased. In contrast, our method is robust to such confounding effects, thus producing more sensible estimates.
@@
The validity of our key identification strategy (along with our estimation procedure) relies only on monotonicity in an index structure, and therefore it has wider applicability beyond panel multinomial choice models. We also introduce the multi-index single-crossing (MISC) condition framework, a general econometric framework under which our method can be applied. This framework encompasses the key ingredients of a large class of models, such as binary choice models with awareness, binary choice with endogeneity, dyadic network formation, bilateral matching, and endogenous censoring.
@@
This paper builds upon and contributes to a large literature on semiparametric (and parametric) discrete choice models, dating back to McFadden (1974) and Manski (1975), and more specifically a recent branch of research that focuses on panel multinomial choice models. Our work is most closely related to the work by Pakes and Porter (2024), who also exploit weak monotonicity and time homogeneity, but restrict the effect of unobserved heterogeneity to be a scalar index that is additively separable from the index of observable characteristics. Another related paper is Shi, Shum, and Song (2018), who exploit cyclical monotonicity of vector-valued functions in a fully additive panel multinomial choice model. Khan, Ouyang, and Tamer (2021) consider another additive model, but utilize the subsample of observations with time-invariant covariates along all products but one so as to leverage univariate monotonicity. Relatedly, the earlier work by Honoré and Kyriazidou (2000) also exploits univariate monotonicity when certain covariates across two periods are equal in a dynamic panel setting. Another recent paper by Chernozhukov, Fernández-Val, and Newey (2019) consider a model with an additive effect under an “on-the-diagonal” restriction (i.e., when covariates at two different time periods coincide). Our method is significantly different from and thus complementary to those proposed in these afore-cited papers.
@@
At a more general level, our work can be related to and compared to semiparametric methods of identification and estimation in models that feature monotonicity in a single parametric index. A related class of estimators that leverage univariate monotonicity, known as maximum score or rank-order estimators, date back to a series of important contributions by Manski (1975, 1985, 1987), and are further investigated in Han (1987), Horowitz (1992), Abrevaya (2000), Honoré and Lewbel (2002) and Fox (2007). Despite the similarity in the reliance on monotonicity, the multinomial or multi-index nature of our current model induces a key difference from the single-index setting, leading to a significantly different method of estimation relative to rank-order estimators.
@@
The rest of this paper is organized as follows. Section 2 introduces our main model specifications and assumptions. Section 2.2 presents our key identification strategy. In Section 3 we provide consistent estimators along with a computational procedure to implement it. Section 4 discusses the generalization of our method to the framework of multi-index single-crossing conditions. Section 5 and Section 6 switch back to our main panel multinomial model, for which we provide a simulation study and an empirical illustration using the Nielsen data.
@@ 2.1 Model and Assumptions
In this section we present a semiparametric panel multinomial choice model featured by infinite-dimensional unobserved heterogeneity and flexible forms of nonseparability, which we will use as the main model to illustrate our identification and estimation method.
Specifically, we consider the following model, which states that individual $i$ chooses product $j$ at time $t$ if and only if $i$ prefers product $j$ to all other alternatives at time $t$:
$y_{ijt}$ $=\mathbf{\mathbbm{1}}\left\{u\left(X_{ijt}^{{}^{\prime}}\beta_{0},\,A_{ij},\,\epsilon_{ijt}\right)>\max_{k\in\left\{1,...,J\right\}\backslash\left\{j\right\}}u\left(X_{ikt}^{{}^{\prime}}\beta_{0},\,A_{ik},\,\epsilon_{ikt}\right)\right\}$ (1)
@@
  • $i\in\left\{1,...N\right\}$ denotes $N$ individuals.
  • $j\in{\cal J}:=\left\{1...,J\right\}$ denotes $J$ choice alternatives, with $J$ products indexed by $1,...,J$. In this paper, we take the number of products $J$ to be fixed.
  • $t\in\left\{1,...,T\right\}$ denotes $T\geq 2$ different time periods. In this paper, we consider a short-panel setting where $T$ is fixed.
  • $X_{ijt}$ is $\mathbb{R}^{d}$-valued vector of observable characteristics specific to each agent-product-time tuple $ijt$. This could include, for example, buyer characteristics such as income level, product characteristics such as price and promotion status, as well as interaction and higher-order terms of those characteristics.
  • $y_{ijt}$ is an observable binary variable, with $y_{ijt}=1$ indicating that buyer $i$ chooses products $j$ at time $t$ and $y_{ijt}=0$ indicating otherwise.
  • $\beta_{0}\in\mathbb{R}^{d}$ is a finite-dimensional unknown parameter of interest.
  • $A_{ij}$ represents an $ij$-specific time-invariant unobserved heterogeneity term of arbitrary dimensions, which we will refer to as the ($ij$-specific) fixed effect.
  • $\epsilon_{ijt}$ is an $ijt$-specific unobserved error term of arbitrary dimensions, which captures time-idiosyncratic utility shocks to product $j$ for agent $i$ at time $t$.
  • $u$ is an unknown function, interpreted as a utility function that aggregates the parametric index $X_{ijt}^{{}^{\prime}}\beta_{0}$, the fixed effect $A_{ij}$ and the error term $\epsilon_{ijt}$ into a scalar representing agent $i$’s utility from choosing product $j$ at time $t$.
Now, we first present our main modeling assumptions, and will follow up with a discussion of these assumptions along with our model specification (1).
To economize on notation, we will from now on frequently refer to the collection of variables concatenated along product and time dimensions: ${\bf X}_{it}:=\left(X_{ijt}\right)_{j=1}^{J}$, ${\bf X}_{i}=\left({\bf X}_{it}\right)_{t=1}^{T}$, ${\bf A}_{i}:=\left(A_{ij}\right)_{j=1}^{J}$, $\boldsymbol{\epsilon}_{it}=\left(\epsilon_{ijt}\right)_{j=1}^{J}$ and $\boldsymbol{\epsilon}_{i}=\left(\boldsymbol{\epsilon}_{it}\right)_{t=1}^{T}$. Recall that we defined $\delta_{ijt}:=X_{ijt}^{{}^{\prime}}\beta_{0}$. The first assumption below imposes a monotonicity restriction on the utility function.
@@
$\left({\bf Y}_{i},{\bf X}_{i},{\bf A}_{i},\boldsymbol{\epsilon}_{i}\right)$ is i.i.d. across $i\in\left\{1,...,N\right\}$ with $N\to\infty$.
Assumption 1 is a standard assumption. [^1] It is worth noting that so far we have not made any explicit restriction on the structure of the spaces on which the arbitrary dimensional random elements ${\bf A}_{i}$ and $\boldsymbol{\epsilon}_{i}$ are defined, but implicit in our specification as well as Assumption 1 is the requirement that $\left({\bf Y}_{i},{\bf X}_{i},{\bf A}_{i},{\bf\epsilon}_{i}\right)$ be well-defined as random elements (measurable functions) on a sufficiently rich probability space $\left(\Omega,\mathscr{F},\mathbb{P}\right)$. Recall that the number of time periods $T$ is held fixed at a finite number, and we focus on a short panel setting with cross-sectional asymptotics.
@@
For every realization of $\left(A_{ij},\epsilon_{ijt}\right)$, the real-valued mapping defined by $\tilde{\delta}\longmapsto u\left(\tilde{\delta},A_{ij},\epsilon_{ijt}\right)$ is weakly increasing in the scalar-valued argument $\tilde{\delta}$.
Essentially, Assumption 2 states that $u\left(X_{ijt}^{{}^{\prime}}\beta_{0},\,A_{ij},\,\epsilon_{ijt}\right)$, individual $i$’s utility of choosing product $j$ at time $t$, is increasing [^2] It should be clarified that increasingness is without loss of generality given monotonicity, since the index $\delta_{ijt}=X_{ijt}^{{}^{\prime}}\beta_{0}$ contains an unknown parameter with unrestricted signs. in the parametric index $X_{ijt}^{{}^{\prime}}\beta_{0}$. Given the parametric index structure $X_{ijt}^{{}^{\prime}}\beta_{0}$, monotonicity itself is a very weak assumption. In the standard panel multinomial choice model with scalar-valued $A_{ij}$ and $\epsilon_{ij}$ along with additive $u$, i.e., $u\left(X_{ijt}^{{}^{\prime}}\beta_{0},\,A_{ij},\,\epsilon_{ijt}\right)=X_{ijt}^{{}^{\prime}}\beta_{0}+A_{ij}+\epsilon_{ijt}$, Assumption 2 is trivially satisfied (with strictness) by construction.
In a way, Assumption 2 endows the parametric index $X_{ijt}^{{}^{\prime}}\beta_{0}$ and the parameter $\beta_{0}$ with economic interpretations. Under Assumption 2, $X_{ijt}^{{}^{\prime}}\beta_{0}$ may be considered as a quality measure of the match between agent $i$ and product $j$ based on their observable characteristics at time $t$, inducing a consequent interpretation of the parameter $\beta_{0}$ as representing how a certain change in a linear combination of observable characteristics may increase utilities for all agents from a certain product $j$, ceteris paribus. Hence, without Assumption 2, it would be hard to interpret $\beta_{0}$.
Moreover, the monotonicity restriction is imposed on the index $X_{ijt}^{{}^{\prime}}\beta_{0}$, but not directly on any specific observable characteristics in $X_{ijt}$: quadratic or higher-order polynomial terms as well as other nonlinear or non-monotone functions of observable characteristics may be included in $X_{ijt}$ whenever appropriate.
@@
The distributions of $\boldsymbol{\epsilon}_{it}$ and $\boldsymbol{\epsilon}_{is}$ conditional on $\left({\bf X}_{is},{\bf X}_{it},{\bf A}_{i}\right)$ are the same across any pair of periods $s\neq t\in\left\{1,...,T\right\}$, i.e.,
$\left.\boldsymbol{\epsilon}_{is}\right|\left({\bf X}_{is},{\bf X}_{it},{\bf A}_{i}\right)\left.\sim\boldsymbol{\epsilon}_{it}\right|\left({\bf X}_{is},{\bf X}_{it},{\bf A}_{i}\right).$

### v3

@@ Abstract
This paper proposes a robust method for semiparametric identification and estimation in panel multinomial choice models, where we allow for infinite-dimensional fixed effects that enter into consumer utilities in an additively nonseparable way, thus incorporating rich forms of unobserved heterogeneity. Our identification strategy exploits multivariate monotonicity in parametric indices, and uses the logical contraposition of an intertemporal inequality on choice probabilities to obtain identifying restrictions. We provide a consistent estimation procedure, and demonstrate the practical advantages of our method with Monte Carlo simulations and an empirical illustration on popcorn sales with the NielsenIQ data.
@@ 1 Introduction
This paper proposes a method for semiparametric identification and estimation in panel multinomial choice models, where we allow for infinite-dimensional fixed effects that enter into consumer utilities in an additively nonseparable manner. The proposed method also applies more widely beyond panel multinomial choice models, and can be adapted to a wide range of models characterized by multi-index single-crossing conditions, which we introduce later in this paper.
@@
$y_{ijt}=\mathbf{\mathbbm{1}}\left\{u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)\geq\max_{k\in\left\{1,\ldots,J\right\}}u\left(X_{ikt}^{{}^{\prime}}\beta_{0},A_{ik},\epsilon_{ikt}\right)\right\},$
@@
where agent $\mathit{i}$’s utility from a candidate product $j$ at time $t$, represented by $u(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt})$, is taken to be a function of three components. The first is a linear index $X_{ijt}^{{}^{\prime}}\beta_{0}$ of observable characteristics $X_{ijt}$, which contains a finite-dimensional parameter of interest $\beta_{0}$ we will identify and estimate. The second term $A_{ij}$ is an infinite-dimensional fixed effect that can be heterogeneous across each agent-product combination. We emphasize that $X_{ijt}$ and $A_{ij}$ can be arbitrarily dependent. The last term $\epsilon_{ijt}$ is an idiosyncratic time-varying error term of arbitrary dimension. The three components are then aggregated by an unknown utility function $u$ in an additively nonseparable way, with the only restriction being that each agent’s utility $u(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt})$ is increasing in its first argument. Each agent then chooses a certain product in a given time period, represented by $y_{ijt}=1$, if and only if this product gives her the highest utility among all available products.
The infinite dimensionality of the terms $u$, $A_{ij}$, and $\epsilon_{ijt}$, together with the model’s additively non-separable interaction structure, jointly generates a rich class of unobserved heterogeneity. Across each agent-product combination $ij$, we are effectively allowing for nonparametric variations in agent utilities. Such variation proxies for the effects of complicated unobserved factors that influence choice behavior, such as brand loyalty, subtle flavors, and unique styles of products. In addition, we work with a nonparametric time homogeneity assumption on the error terms $\epsilon_{ijt}$ that restricts $\epsilon_{ijt}$ and $\epsilon_{ijs}$ from two periods $t$ and $s$ to have the same marginal distribution given the observed covariates from the two periods. Apart from this, we impose no parametric restrictions on the distribution of $\epsilon_{ijt}$ and no additional restrictions on its dependence across time $t$ and products $j$. In particular, the fully unrestricted dependence of $\epsilon_{ijt}$ across products $j$ allows our framework to remain robust to the well-known “Blue-Bus/Red-Bus” problem and related pathologies that arise in many standard multinomial choice models, as discussed in Berry and Pakes (2007). [^1] See the discussion after Assumption 3 in Section 2.1 for more details.
@@
The generality of our framework nests many semiparametric (and parametric) panel multinomial choice models with scalar fixed effects, scalar error terms, and varying degrees of additive separability, including the following standard specification:
@@
$y_{ijt}=\mathbf{\mathbbm{1}}\left\{X_{ijt}^{{}^{\prime}}\beta_{0}+A_{ij}+\epsilon_{ijt}>\max_{k\in\left\{1,\ldots,J\right\}}\left(X_{ikt}^{{}^{\prime}}\beta_{0}+A_{ik}+\epsilon_{ikt}\right)\right\}.$
@@
Relative to existing work, our framework accommodates both the infinite dimensionality of unobserved heterogeneity and non–additive separability in agent utilities, under a standard time-homogeneity assumption on the idiosyncratic error term that is widely used in the literature.
@@
Our identification strategy leverages multivariate monotonicity in contrapositive form. The intuition is straightforward: if the choice probability of a given product (or subset of products) strictly increases from one period to the next, then it cannot be that this product (or all products in the subset) becomes worse while all other products become better over the two periods. By applying this contraposition to a carefully constructed inequality in conditional choice probabilities, we obtain an identifying restriction on the index values that is free of all infinite-dimensional nuisance parameters. We further show that, in a two-period setting, the identified set obtained by aggregating these restrictions across all product subsets is sharp.
@@
Based on our identification result, we provide consistent two-step set (or point) estimators, together with a computational algorithm adapted to the technical challenges of our framework. The first stage takes the form of a standard nonparametric regression, where we estimate a collection of intertemporal differences in conditional choice probabilities. In the second stage, we numerically minimize our sample criterion function with the first-stage estimates plugged in. A highlight of our computational procedure is the adoption of a spherical-coordinate reparameterization of our criterion functions in terms of angles, which enables us to exploit a combination of topological, geometric and computational advantages. A simulation study is conducted to analyze the finite-sample performance of our method and the adequacy of our computational procedure for practical implementation.
@@
We also provide an empirical illustration of our procedure, where we use the NielsenIQ data on popcorn sales in the United States to explore the effects of marketing promotion. The results show that our procedure produces estimates that conform well with economic intuition. For example, we find that special in-store displays boost sales not only through a direct promotion effect but also through the attenuation of consumer price sensitivity. Intuitively, marketing managers are more likely to promote products for which they know consumers are more sensitive to price and promotion. Hence, the average effective price sensitivities of promoted products tend to be larger than those not promoted due to the selection effect. Given the non-additive nature of such selection effects, estimators based on additive separability will be biased. In contrast, our method is robust to such confounding effects, thus producing more sensible estimates.
@@
The validity of our identification strategy, as well as our estimation procedure, relies solely on monotonicity in an index structure, and thus extends naturally beyond panel multinomial choice models. We also introduce the multi-index single-crossing (MISC) condition framework, a general econometric framework under which our method can be applied. This framework encompasses the key ingredients of a large class of models, such as binary choice models with awareness, binary choice with endogeneity, dyadic network formation, bilateral matching, and endogenous censoring.
We acknowledge several limitations of our identification approach. First, our current model setup does not allow for time-varying endogeneity between observed covariates $X_{ijt}$ and the error term $\epsilon_{ijt}$, which effectively rules out the inclusion of contemporaneously endogenous covariates and/or lagged outcomes. See a subsequent paper by Gao and Wang (2025) for a weaker version of the time homogeneity assumption that can be exploited for identification in panel multinomial choice models with endogenous and dynamic covariates. Second, our current approach does not allow for random coefficients on time-varying covariates. That said, rich forms of time-invariant taste heterogeneity have been absorbed into the nonparametric fixed effects under our current setting. Third, in the short-panel setting, our current approach effectively “differences out” the unobserved individual fixed effects $A_{i}$. As a result, it cannot identify counterfactual parameters that depend on the distribution of $A_{i}$. However, in long panels, our approach can be adapted to identify counterfactual parameters, which we discuss in more detail in Appendix E.
@@
This paper builds upon and contributes to a large literature on semiparametric (and parametric) discrete choice models, dating back to McFadden (1974) and Manski (1975), and more specifically to the line of literature on panel multinomial choice models. Our work is most closely related to Pakes and Porter (2024), who also exploit weak monotonicity and time homogeneity, but restrict the effect of unobserved heterogeneity to be a scalar index that is additively separable from the index of observable characteristics. Shi, Shum, and Song (2018) exploit cyclical monotonicity of vector-valued functions in a fully additive panel multinomial choice model. Khan, Ouyang, and Tamer (2021) consider another additive model, but utilize the subsample of observations with time-invariant covariates along all products but one so as to leverage univariate monotonicity. Honoré and Kyriazidou (2000) also exploit univariate monotonicity when certain covariates across two periods are equal in a dynamic panel setting. Chernozhukov, Fernández-Val, and Newey (2019) consider a model with an additive effect under an “on-the-diagonal” restriction (i.e., when covariates at two different time periods coincide). By allowing non-additiveness in the specification of utility functions and infinite-dimensional fixed effects, our method is different from and thus complementary to those proposed in these aforementioned papers.
This paper is also connected to the literature on the nonparametric identification and estimation in discrete choice models (e.g., Berry and Haile (2014); Compiani (2022)). That literature assumes monotonicity restrictions of the demand functions in product-market specific parametric indices to invert the demand system. This is particularly useful as it leads to a system of equations with only one unobservable per equation, from which the unobservable product-market specific demand shifter can be successfully constructed. Our paper considers a different model, but also leverages monotonicity restrictions in parametric indices to facilitate identification and estimation of the structural parameters. In both cases, the index assumption restricts the amount of unobserved heterogeneity in preferences for the variables included in the index.
@@
More broadly, our work connects to the semiparametric literature on the identification and estimation of models characterized by monotonicity in a single parametric index. A related class of estimators that leverage univariate monotonicity, known as maximum score or rank-order estimators, dates back to a series of important contributions by Manski (1975, 1985, 1987), and is further investigated in Han (1987), Horowitz (1992), Abrevaya (2000), Honoré and Lewbel (2002), Fox (2007), and Yan and Yoo (2019). [^2] We clarify that Manski (1975), Fox (2007) and Yan and Yoo (2019) consider multinomial choice models with multiple parametric indices, but focus on settings where the “multi-index” problem can be reduced to leverage single-variate monotonicity (or a single-variate rank-order property). Despite the similar reliance on monotonicity, the multi-index nature of our model, and more importantly the multivariate monotonicity condition that we leverage, induce key differences from the single-index (and univariate monotonicity) setting, leading to a substantially different estimation method relative to rank-order estimators.
@@
The rest of this paper is organized as follows. Section 2 introduces our main model specifications and assumptions. Section 2.2 presents our key identification strategy. In Section 3, we provide consistent estimators along with a computational procedure to implement it. Section 4 discusses the generalization of our method to the framework of multi-index single-crossing conditions. Sections 5 and 6 switch back to our main panel multinomial choice model, for which we provide a simulation study and an empirical illustration using the NielsenIQ data. We conclude in Section 7.
@@ 2.1 Model and Assumptions
In this section, we present a semiparametric panel multinomial choice model featuring infinite-dimensional unobserved heterogeneity and flexible forms of non-separability, which serves as the main framework for illustrating our identification and estimation method.
Specifically, we consider the following model, in which individual $i$ chooses product $j$ at time $t$ if and only if $i$ prefers product $j$ to all other alternatives at time $t$:
[…]
@@
  • $i\in\{1,\ldots,N\}$ denotes $N$ individuals.
  • $j\in{\cal J}:=\{1,\ldots,J\}$ denotes the set of $J$ choice alternatives, with products indexed by $1,\ldots,J$. Throughout this paper, we treat the number of products $J$ as fixed.
  • $t\in\{1,\ldots,T\}$ denotes the $T\geq 2$ time periods. In this paper, we consider a short-panel setting in which $T$ is fixed.
  • $X_{ijt}$ is an $\mathbb{R}^{D}$-valued vector of observable characteristics specific to each agent–product–time tuple ijtijt. These may include, for example, buyer characteristics such as income, product characteristics such as price and promotion status, as well as interaction and higher-order terms of these variables.
  • $y_{ijt}$ is an observable binary variable, with $y_{ijt}=1$ indicating that buyer $i$ chooses product $j$ at time $t$, and $y_{ijt}=0$ indicating otherwise.
  • $\beta_{0}\in\mathbb{R}^{D}$ is the finite-dimensional parameter of interest.
  • $A_{ij}$ represents an $ij$-specific time-invariant unobserved heterogeneity term of arbitrary dimension, which we refer to as the $ij$-specific fixed effect.
  • $\epsilon_{ijt}$ is an ijtijt-specific unobserved error term of arbitrary dimension, which captures time-idiosyncratic utility shocks to product $j$ for agent $i$ at time $t$.
  • $u$ is an unknown function, interpreted as a utility function that aggregates the parametric index $X_{ijt}^{{}^{\prime}}\beta_{0}$, the fixed effect $A_{ij}$, and the error term $\epsilon_{ijt}$ into a scalar representing agent $i$’s utility from choosing product $j$ at time $t$.
We first present our main modeling assumptions and then discuss these assumptions in conjunction with our model specification (1). To economize on notation, we refer to the collection of variables concatenated along product and time dimensions: ${\bf X}_{it}=(X_{ijt})_{j=1}^{J}$, ${\bf X}_{i}=({\bf X}_{it})_{t=1}^{T}$, ${\bf A}_{i}=(A_{ij})_{j=1}^{J}$, $\boldsymbol{\epsilon}_{it}=(\epsilon_{ijt})_{j=1}^{J}$, and $\boldsymbol{\epsilon}_{i}=(\boldsymbol{\epsilon}_{it})_{t=1}^{T}$. We also write $\delta_{ijt}=X_{ijt}^{{}^{\prime}}\beta_{0}$ to denote the parametric index.
@@
$({\bf Y}_{i},{\bf X}_{i},{\bf A}_{i},\boldsymbol{\epsilon}_{i})$ is i.i.d. across $i\in\left\{1,\ldots,N\right\}$ with $N\to\infty$.
Assumption 1 is a standard assumption. [^3] It is worth noting that we have not yet imposed any explicit restrictions on the structure of the spaces in which the arbitrary-dimensional random elements ${\bf A}_{i}$ and $\boldsymbol{\epsilon}_{i}$ are defined. However, implicit in our model specification and in Assumption 1 is the requirement that $({\bf Y}_{i},{\bf X}_{i},{\bf A}_{i},{\bf\epsilon}_{i})$ be well-defined as random elements—that is, measurable functions—on a sufficiently rich probability space $(\Omega,\mathscr{F},\mathbb{P})$. Recall that the number of time periods $T$ is held fixed, and we focus on a short panel setting with cross-sectional asymptotics.
@@
For every realization of $(A_{ij},\epsilon_{ijt})$, the mapping $\tilde{\delta}\longmapsto u(\tilde{\delta},A_{ij},\epsilon_{ijt})$ is weakly increasing in the scalar-valued argument $\tilde{\delta}$.
Essentially, Assumption 2 states that $u(\delta_{ijt},A_{ij},\epsilon_{ijt})$, the utility of individual $i$ from choosing product $j$ at time $t$, is weakly increasing [^4] It should be clarified that increasingness is without loss of generality given monotonicity, since the index $\delta_{ijt}=X_{ijt}^{{}^{\prime}}\beta_{0}$ contains an unknown parameter with unrestricted signs. in the index $\delta_{ijt}$. Given the index structure, monotonicity itself is a relatively mild assumption. In the standard panel multinomial choice model with scalar-valued $A_{ij}$ and $\epsilon_{ij}$ along with additive $u$, i.e., $u(\delta_{ijt},A_{ij},\epsilon_{ijt})=\delta_{ijt}+A_{ij}+\epsilon_{ijt}$, Assumption 2 is trivially satisfied (with strictness) by construction.
In a way, Assumption 2 endows the index $\delta_{ijt}$ and the parameter $\beta_{0}$ with economic interpretations. Under Assumption 2, $\delta_{ijt}$ may be considered as a quality measure of the match between agent $i$ and product $j$ based on their observable characteristics at time $t$, inducing an interpretation of $\beta_{0}$ as representing how a certain change in a linear combination of observable characteristics may increase utilities for all agents from a certain product $j$, ceteris paribus. Hence, without Assumption 2, it would be hard to interpret $\beta_{0}$.
Moreover, the monotonicity restriction is imposed on $\delta_{ijt}$, but not directly on any specific observable characteristics in $X_{ijt}$: quadratic or higher-order polynomial terms, and other functions of observable characteristics can be included in $X_{ijt}$ whenever appropriate.
@@
The distributions of $\boldsymbol{\epsilon}_{it}$ and $\boldsymbol{\epsilon}_{is}$ conditional on $({\bf X}_{it},{\bf X}_{is},{\bf A}_{i})$ across any pair of periods $t\neq s\in\{1,\ldots,T\}$ satisfy
$\left.\boldsymbol{\epsilon}_{it}\right|\left({\bf X}_{it},{\bf X}_{is},{\bf A}_{i}\right)\left.\sim\boldsymbol{\epsilon}_{is}\right|\left({\bf X}_{it},{\bf X}_{is},{\bf A}_{i}\right).$